OpenStudy (anonymous):

For what values of k such that kx^2 -10x +5k=0, k NOT=0, will have 2 real roots?

OpenStudy (anonymous):

you want the discriminant to be positive, so solve for \(k\) \[100-4\times k\times 5k>0\]

OpenStudy (anonymous):

i.e. \[b^2-4ac>0\] with \(a=k,b=-10,c=5k\)

OpenStudy (mathstudent55):

A quadratic equation ax^2 + bx + c = 0 has two roots when the discriminant of the quadratic formula is positive. The discriminant is b^2 - 4ac. In your case, a = k, b = -10, and c = 5k. Plug those values into the discriminant and set it greater than zero and solve the inequality.

OpenStudy (anonymous):

... where'd you get that formula? satelite?

OpenStudy (mathstudent55):

It's what's inside the radical of the quadratic formula.

OpenStudy (anonymous):

that's what I was forgeting! discriminate! I kept solving it wrong! thanks!

OpenStudy (anonymous):

it is the discriminant, that is \(b^2-4ac\) from the term inside the radical in the quadratic formula is it clear how to solve \[100-20k^2>0\]?

OpenStudy (anonymous):

Sorry, I'm back... So I now have  But where do I go from here?